# Chapter 4 Discrete subgroups of semisimple Lie groups, beyond lattices

### 26 Feb 2024 v1

@@ 5.2.1 Two attempts
Suppose that $\Gamma_{0}$ acts convex cocompactly via $\rho$ on some properly convex open subset $\Omega$ of $\mathbb{P}(\mathbb{R}^{d})$ such that $\partial\Omega$ is $C^{1}$ and contains no segments. Because $\partial\Omega$ contains no segments, the geodesic rays of $\Omega$ for the Hilbert metric $\mathtt{d}_{\Omega}$ are exactly the projective segments between a point of $\Omega$ and a point of $\partial\Omega$, and two such rays remain at bounded Hausdorff distance for $\mathtt{d}_{\Omega}$ if and only if their endpoints in $\partial\Omega$ are the same. By convexity, the convex core $\mathcal{C}^{\mathsf{cor}}_{\rho(\Gamma_{0})}$, endowed with the restriction of $\mathtt{d}_{\Omega}$, is a geodesic metric space whose visual boundary $\partial_{\infty}\mathcal{C}^{\mathsf{cor}}_{\rho(\Gamma_{0})}$ identifies with its ideal boundary $\overline{\mathcal{C}^{\mathsf{cor}}_{\rho(\Gamma_{0})}}\cap\partial\Omega$ in $\partial\Omega$.
@@
Using the fact that $\partial\Omega$ contains no segments, one can check by a limiting argument that all triangles in $\mathcal{C}^{\mathsf{cor}}_{\rho(\Gamma_{0})}$ must be uniformly thin, i.e. the metric space $(\mathcal{C}^{\mathsf{cor}}_{\rho(\Gamma_{0})},\mathtt{d}_{\Omega})$ is Gromov hyperbolic: see [53, Lem. 6.3]. Since the action of $\Gamma_{0}$ on $(\mathcal{C}^{\mathsf{cor}}_{\rho(\Gamma_{0})},\mathtt{d}_{\Omega})$ via $\rho$ is properly discontinuous, by isometries, with compact quotient, we deduce that $\Gamma_{0}$ is Gromov hyperbolic and (as in the proof of (1$\Rightarrow$ (5) in Section 4.4) that any orbital map $\Gamma_{0}\to\mathcal{C}^{\mathsf{cor}}_{\rho(\Gamma_{0})}$ extends to a continuous $\rho$-equivariant boundary map $\xi_{1}:\partial_{\infty}\Gamma_{0}\to\partial_{\infty}\mathcal{C}^{\mathsf{cor}}_{\rho(\Gamma_{0})}\subset\mathbb{P}(\mathbb{R}^{d})$.
@@
Consider the dual convex set $\Omega^{*}=\{H\in\mathbb{P}((\mathbb{R}^{d})^{*})\,|\,H\cap\overline{\Omega}=\emptyset\}$ (where we view $\mathbb{P}((\mathbb{R}^{d})^{*})$ as the set of projective hyperplanes in $\mathbb{P}(\mathbb{R}^{d})$). The boundary $\partial\Omega^{*}$ of $\Omega^{*}$ is $C^{1}$ (because $\partial\Omega$ contains no segments), and it contains no segments (because $\partial\Omega$ is $C^{1}$). One can show that the dual action of $\Gamma_{0}$ on $\Omega^{*}$ via $\rho$ is still convex cocompact: see [53, § 5]. Then the same reasoning as above yields a continuous $\rho$-equivariant boundary map $\xi_{d-1}:\partial_{\infty}\Gamma_{0}\to\mathbb{P}((\mathbb{R}^{d})^{*})$.
@@
By construction, $\xi_{1}$ and $\xi_{d-1}$ are transverse: indeed, $\xi_{d-1}(w)$ is a supporting hyperplane to $\Omega$ at $\xi_{1}(w)$, and $\partial\Omega$ contains no segments. To conclude, we show that $\xi_{1}$ and $\xi_{d-1}$ are dynamics-preserving and (using Lemma 4.40) that $(\mu_{1}-\mu_{2})(\rho(\gamma))\to+\infty$ as $\mathtt{d}_{\mathrm{Cay}(\Gamma_{0})}(e,\gamma)\to+\infty$: see [53, § 7]. We then apply the implication (4$\Rightarrow$ (1) of Theorem 4.37. ∎
5.2.4 Applications
Theorem 4.44 and Corollaries 4.454.46 give geometric interpretations for Anosov representations.
@@ 5.2.4 Applications
For odd $d$, any Hitchin representation $\rho:\pi_{1}(S)\to\mathrm{PSL}(d,\mathbb{R})$ as in Section 3.3 preserves a nonempty properly convex open subset of $\mathbb{P}(\mathbb{R}^{d})$ (see [52, 53, 139]). Therefore these representations are strongly convex cocompact in $\mathbb{P}(\mathbb{R}^{d})$ by Theorem 4.44. This extends the case $d=3$ due to Choi and Goldman (see the proof of Theorem 4.11).
@@
For $n\geq 2$, any maximal representation $\rho:\pi_{1}(S)\to\mathrm{SO}(2,n)$ as in Section 3.3 preserves a nonempty properly convex open subset of $\mathbb{P}(\mathbb{R}^{n+2})$, contained in $\mathbb{H}^{2,n-1}=\{[v]\in\mathbb{P}(\mathbb{R}^{n+2})\,|\,\langle v,v\rangle_{2,n}<\nolinebreak 0\}$ (see [46, 52]). Therefore these representations are strongly convex cocompact in $\mathbb{P}(\mathbb{R}^{n+2})$ by Theorem 4.44, and in fact $\mathbb{H}^{2,n-1}$-convex cocompact as in Section 3.4 (see the comments after Theorem 4.44).
@@
where $m_{i,i}=1$ (i.e. $s_{i}$ is an involution) and $m_{i,j}\in\{2,3,4,\dots\}\cup\{\infty\}$ for all $i\neq j$. (By convention, $(s_{i}s_{j})^{\infty}=e$ means that $s_{i}s_{j}$ has infinite order in the group $W$.) Vinberg [128] developed a theory of representations of $W$ as a reflection group in a finite-dimensional real vector space $V$: these are by definition representations $\rho:W\to\mathrm{GL}(V)$ such that each $\rho(s_{i})$ is a linear reflection in a hyperplane of $V$ and the configuration of these reflections is such that $\rho$ is injective, discrete, and the associated fundamental polytope has nonempty interior. These representations may preserve a nondegenerate quadratic form on $V$ (e.g. the image of $\rho$ could be a discrete subgroup of $\mathrm{O}(n,1)$ generated by orthogonal reflections in the faces of a right-angled polyhedron of $\mathbb{H}^{n}$ as in Section 4.3), but in general they need not preserve any nonzero quadratic form. Representations of $W$ as a reflection group constitute a subset $\mathrm{Hom}_{\mathrm{refl}}(W,\mathrm{GL}(V))$ of $\mathrm{Hom}(W,\mathrm{GL}(V))$ which is semialgebraic (defined by finitely many equalities and inequalities).
@@
Let $W$ be a Coxeter group in $N$ generators as in (7). Suppose $W$ is infinite and Gromov hyperbolic. Then for any $d\geq N$ there exist representations $\rho:W\to\mathrm{SL}^{\pm}(d,\mathbb{R})$ of $W$ as a reflection group which are strongly convex cocompact in $\mathbb{P}(\mathbb{R}^{d})$; they constitute (at least for $d\geq 2N-2$) the full interior of $\mathrm{Hom}_{\mathrm{refl}}(W,\mathrm{GL}(d,\mathbb{R}))$. By Theorem 4.44, these representations are $P_{1}$-Anosov.
@@
One can also use the geometric interpretation of Anosov representations from Theorem 4.44 to prove that free products $\Gamma_{1}*\Gamma_{2}$ of Anosov subgroups $\Gamma_{1},\Gamma_{2}$ are Anosov, using a generalisation of the ping pong arguments of Sections 3.13.2. For instance, for $11\leq i\leq d-$ 1, let $\tau_{i}:\mathrm{SL}(d,\mathbb{R})\to\mathrm{SL}(S^{2}(\Lambda^{i}\mathbb{R}^{d}))$ be the second symmetric power of the $i$-th exterior power of the standard representation as in Corollary 4.46, let $V^{\prime}_{i}:=S^{2}(\Lambda^{i}\mathbb{R}^{d})\oplus\mathbb{R}$, and let $\tau^{\prime}_{i}:\mathrm{SL}(d,\mathbb{R})\to\mathrm{SL}(V^{\prime}_{i})$ be the direct sum of $\tau_{i}$ and of the trivial representation. Then the following holds.
@@
5.3 Generalisations of Anosov subgroups
In the past few years, several fruitful generalisations of Anosov subgroups have appeared, which are currently being actively investigated. These generalisations exploit both the dynamical definition of Anosov subgroups from Section 5.1 and their geometric characterisation from Section 5.2. Let us briefly mention three of these generalisations.
5.3.1 More general convex cocompact subgroups
We just saw in Theorem 4.44 and Corollaries 4.45 and 4.46 that Anosov representations can be characterised geometrically by a strong convex cocompactness condition in projective space. Here strong refers to the regularity imposed on the properly convex open set $\Omega$ (its boundary $\partial\Omega$ should be $C^{1}$ and contain no segments).
@@ 5.3.1 More general convex cocompact subgroups
As a generalisation of Example 4.49, let $W$ be a Coxeter group in $N$ generators as in (7). Suppose $W$ is infinite. Then there exists a representation $\rho\in\mathrm{Hom}_{\mathrm{refl}}(W,\mathrm{GL}(d,\mathbb{R}))$ which is convex cocompact in $\mathbb{P}(\mathbb{R}^{d})$ for some $d$ if and only if any affine standard subgroup of $W$ is of type $\tilde{A}_{k}$ for some $k\geq 1$ and $W$ does not contain a direct product of two infinite standard subgroups. If this holds, then we can take any $d\geq N$ and the convex cocompact representations then constitute a large open subset of $\mathrm{Hom}_{\mathrm{refl}}(W,\mathrm{GL}(d,\mathbb{R}))$: see [55, § 1.5].
Examples 4.53, 4.54, and 4.55 provide many convex cocompact groups which are not Gromov hyperbolic. (In Example 4.55, the group $W$ is nonhyperbolic as soon as it contains an affine standard subgroup of type $\tilde{A}_{k}$ with $k\geq 2$, see Remark 4.22.)
Some of these groups are still relatively hyperbolic: e.g. in Example 4.55, the group $W$ is relatively hyperbolic with respect to a collection of virtually abelian subgroups of rank $\geq 2$ (see [55, Cor. 1.7]). We refer to [91] for general results about the structure of relatively hyperbolic groups which are convex cocompact in $\mathbb{P}(\mathbb{R}^{d})$ and about the geometry of the associated convex sets. On the other hand, Example 4.53 includes, for $d=d^{\prime}(d^{\prime}+1)/2\geq 6$, discrete subgroups of $\mathrm{PGL}(d,\mathbb{R})$ which divide a symmetric properly convex open set $\Omega_{\mathrm{sym}}\subset\mathbb{P}(\mathbb{R}^{d})\simeq\mathbb{P}(\mathrm{Sym}_{d^{\prime}}(\mathbb{R}))$ and which are isomorphic to cocompact lattices of $\mathrm{PGL}(d^{\prime},\mathbb{R})$, hence not relatively hyperbolic (see Section 2.3).
In the case that $\Gamma_{0}$ is relatively hyperbolic, Weisman [129] has recently given a dynamical characterisation of convex cocompact representations of $\Gamma_{0}$ that extends the characterisation of Anosov representations of Theorem 4.37.(6). The expansion now takes place in various Grassmannians (not only projective space): namely, at each face of the full orbital limit set in $\partial\Omega$, there is expansion in the Grassmannian of $i$-planes of $\mathbb{R}^{d}$ where $i-1$ is the dimension of the face.
5.3.2 Relatively Anosov subgroups
Kapovich–Leeb [95] and Zhu [135, 136] have developed notions of a relatively Anosov representation of a relatively hyperbolic group into a noncompact semisimple Lie group $G$. They obtain various characterisations similar to those of Theorem 4.37. The original definition of Anosov representations using flows (Definition 4.25 and Condition 4.28) is recovered in this more general setting by recent work of Zhu and Zimmer [137].
Extending Fact 4.30, if $\Gamma_{0}$ is relatively hyperbolic with respect to a collection of subgroups (called peripheral subgroups), then relatively Anosov representations of $\Gamma_{0}$ into a given $G$ are stable under small deformations that preserve the conjugacy class of the image of each peripheral subgroup [95, 137].
@@ 5.3.2 Relatively Anosov subgroups
  • some of the Schottky groups of Section 3.2,
  • the images of certain compositions $\tau\circ\sigma_{0}:\Gamma_{0}\to G$ where $\sigma_{0}:\Gamma_{0}\to G^{\prime}$ is a geometrically finite representation into a semisimple Lie group $G^{\prime}$ with $\mathrm{rank}_{\mathbb{R}}(G^{\prime})=1$ and $\tau:G^{\prime}\to G$ is a representation with compact kernel (e.g. Fact 4.31 generalises to the relative setting);
  • similarly to Section 3.3, small deformations in $G$ of such $\tau\circ\sigma_{0}(\Gamma_{0})$, preserving the conjugacy class of the image of each peripheral subgroup;
  • certain representations of $\mathrm{PSL}(2,\mathbb{Z})$ into $\mathrm{PGL}(3,\mathbb{R})$ constructed by Schwartz [121] by iterating Pappus’s theorem (see [138, § 13]);
  • for a finite-volume hyperbolic surface $S$, the images of positive (in the sense of Fock–Goncharov [69]) type-preserving representations of $\Gamma_{0}=\pi_{1}(S)$ into a real split simple Lie group $G$, see [42] (for closed $S$, these coincide with the Hitchin representations of Sections 3.33.4);
@@
It would be interesting to determine whether relatively Anosov representations of relatively hyperbolic groups can also be fully characterised geometrically similarly to Theorem 4.44 and Corollaries 4.454.46.
5.3.3 Extended geometrically finite subgroups
@@
Recently, Weisman [130, 131] has introduced a notion of extended geometrically finite (or EGF for short) representation of a relatively hyperbolic group. This is a dynamical notion, which extends a dynamical characterisation of Anosov representations in terms of multicones [26]. EGF representations include all Anosov or relatively Anosov representations, all representations of relatively hyperbolic groups which are convex cocompact in the sense of Definition 4.52, as well as other examples (see [131, Th. 1.5–1.7] and [23, Prop. 6.5 & Rem. 6.2]). They are stable under certain small deformations, called peripherally stable, for which the dynamics of the peripheral subgroups does not degenerate too much.
@@
On the other hand, it would be interesting to define a general notion of geometric finiteness in convex projective geometry (involving properly convex open subsets $\Omega$ of $\mathbb{P}(\mathbb{R}^{d})$ where $\partial\Omega$ may contain segments or not be $C^{1}$), and to make the link with Weisman’s EGF representations. A good notion of geometric finiteness should contain as a particular case the notion of convex cocompactness from Definition 4.51. More precisely, a convex projective manifold $M=\Gamma\backslash\Omega$ should be geometrically finite if its convex core $\Gamma\backslash\mathcal{C}^{\mathsf{cor}}_{\Gamma}(\Omega)$ (see Definition 4.51) is covered by a compact piece and finitely many ends of $M$, called cusps, with a controlled geometry. It is not completely clear what the right definition of a cusp should be. Following Cooper, Long, and Tillmann [47], one could define a (full) cusp to be the image in $M$ of some convex open subset of $\Omega$ whose stabiliser in $\Gamma$ is infinite and does not contain any hyperbolic element (i.e. any element of this stabiliser has all its complex eigenvalues of the same modulus); in that case, the cusp is diffeomorphic to the direct product of $\mathbb{R}$ with an affine $(d-2)$-dimensional manifold called the cusp cross-section, and the stabiliser of the cusp is virtually nilpotent [47, Th. 5.3]. The cusp is said to have maximal rank if the cross-section is compact. A more general notion of cusp of maximal rank, where the stabiliser may contain hyperbolic elements but is still assumed to be virtually nilpotent, was studied in [5]. A notion of geometric finiteness involving only such generalised cusps of maximal rank was introduced and studied in [134], where it was characterised in dynamical terms. Examples (both of finite and infinite volume) were constructed in [6, 24] as small deformations of finite-volume real hyperbolic manifolds, using a stability result from [48]; the corresponding representations are EGF by [131]. On the other hand, the study of convex projective cusps of nonmaximal rank, possibly allowing for hyperbolic elements, is still at its infancy, and a good general notion of geometric finiteness in this setting still remains to be found, together with appropriate dynamical characterisations.
Acknowledgements
@@
These notes grew out of a minicourse given at Groups St Andrews 2022 in Newcastle. I would like to warmly thank the organisers (Colin Campbell, Martyn Quick, Edmund Robertson, Colva Roney-Dougal, and David Stewart) for a very stimulating and enjoyable meeting.

### 28 Mar 2024 v2

@@ 2 Anosov representations and convex cocompactness
Suppose that $\Gamma_{0}$ acts convex cocompactly via $\rho$ on some properly convex open subset $\Omega$ of $\mathbb{P}(\mathbb{R}^{d})$ such that $\partial\Omega$ is $C^{1}$ and contains no segments. Because $\partial\Omega$ contains no segments, the geodesic rays of $\Omega$ for the Hilbert metric $\mathtt{d}_{\Omega}$ (Figure 11) are exactly the projective segments between a point of $\Omega$ and a point of $\partial\Omega$, and two such rays remain at bounded Hausdorff distance for $\mathtt{d}_{\Omega}$ if and only if their endpoints in $\partial\Omega$ are the same. Therefore the convex core $\mathcal{C}^{\mathsf{cor}}_{\rho(\Gamma_{0})}$, endowed with the restriction of $\mathtt{d}_{\Omega}$, is a geodesic metric space whose visual boundary $\partial_{\infty}\mathcal{C}^{\mathsf{cor}}_{\rho(\Gamma_{0})}$ identifies with its ideal boundary $\overline{\mathcal{C}^{\mathsf{cor}}_{\rho(\Gamma_{0})}}\cap\partial\Omega$ in $\partial\Omega$.
@@
Using the fact that $\partial\Omega$ contains no segments, one can check by a limiting argument that all triangles in $\mathcal{C}^{\mathsf{cor}}_{\rho(\Gamma_{0})}$ must be uniformly thin, i.e. the metric space $(\mathcal{C}^{\mathsf{cor}}_{\rho(\Gamma_{0})},\mathtt{d}_{\Omega})$ is Gromov hyperbolic: see [54, Lem. 6.3]. Since the action of $\Gamma_{0}$ on $(\mathcal{C}^{\mathsf{cor}}_{\rho(\Gamma_{0})},\mathtt{d}_{\Omega})$ via $\rho$ is properly discontinuous, by isometries, with compact quotient, we deduce that $\Gamma_{0}$ is Gromov hyperbolic and (as in the proof of (1$\Rightarrow$ (5) in Section 4) that any orbital map $\Gamma_{0}\to\mathcal{C}^{\mathsf{cor}}_{\rho(\Gamma_{0})}$ extends to a continuous $\rho$-equivariant boundary map $\xi_{1}:\partial_{\infty}\Gamma_{0}\to\partial_{\infty}\mathcal{C}^{\mathsf{cor}}_{\rho(\Gamma_{0})}\subset\mathbb{P}(\mathbb{R}^{d})$.
@@
Consider the dual $\Omega^{*}=\{H\in\mathbb{P}((\mathbb{R}^{d})^{*})\,|\,H\cap\overline{\Omega}=\emptyset\}$ of $\Omega$ (where we view $\mathbb{P}((\mathbb{R}^{d})^{*})$ as the set of projective hyperplanes in $\mathbb{P}(\mathbb{R}^{d})$). It is a properly convex open subset of $\mathbb{P}((\mathbb{R}^{d})^{*})$. The boundary $\partial\Omega^{*}$ of $\Omega^{*}$ is $C^{1}$ (because $\partial\Omega$ contains no segments), and it contains no segments (because $\partial\Omega$ is $C^{1}$). One can show that the dual action of $\Gamma_{0}$ on $\Omega^{*}$ via $\rho$ is still convex cocompact: see [54, § 5]. Then the same reasoning as above yields a continuous $\rho$-equivariant boundary map $\xi_{d-1}:\partial_{\infty}\Gamma_{0}\to\mathbb{P}((\mathbb{R}^{d})^{*})$.
@@
By construction, $\xi_{1}$ and $\xi_{d-1}$ are transverse: indeed, $\xi_{d-1}(w)$ is a supporting hyperplane to $\Omega$ at $\xi_{1}(w)$ for any $w$, and $\partial\Omega$ contains no segments. One checks that $\xi_{1}$ and $\xi_{d-1}$ are dynamics-preserving and (using Lemma 5.16) that $(\mu_{1}-\mu_{2})(\rho(\gamma))\to+\infty$ as $\mathtt{d}_{\mathrm{Cay}(\Gamma_{0})}(e,\gamma)\to+\infty$: see [54, § 7]. We then apply the implication (4$\Rightarrow$ (1) of Theorem 5.13. ∎
Applications
Theorem 5.20 and Corollaries 5.215.22 give geometric interpretations for Anosov representations.
@@
For odd $d$, any Hitchin representation $\rho:\pi_{1}(S)\to\mathrm{PSL}(d,\mathbb{R})$ as in Section 3 preserves a nonempty properly convex open subset of $\mathbb{P}(\mathbb{R}^{d})$ (see [53, 54, 141]). Therefore these representations are strongly convex cocompact in $\mathbb{P}(\mathbb{R}^{d})$ by Theorem 5.20. This extends the case $d=3$ due to Choi and Goldman (see the proof of Theorem 3.7).
@@
For $n\geq 2$, any maximal representation $\rho:\pi_{1}(S)\to\mathrm{SO}(2,n)$ as in Section 3 preserves a nonempty properly convex open subset of $\mathbb{P}(\mathbb{R}^{n+2})$, contained in $\mathbb{H}^{2,n-1}=\{[v]\in\mathbb{P}(\mathbb{R}^{n+2})\,|\,\langle v,v\rangle_{2,n}<\nolinebreak 0\}$ (see [47, 53]). Therefore these representations are strongly convex cocompact in $\mathbb{P}(\mathbb{R}^{n+2})$ by Theorem 5.20, and in fact $\mathbb{H}^{2,n-1}$-convex cocompact as in Section 4 (see the comments after Theorem 5.20).
@@
where $m_{i,i}=1$ (i.e. $s_{i}$ is an involution) and $m_{i,j}\in\{2,3,4,\dots\}\cup\{\infty\}$ for all $i\neq j$. (By convention, $(s_{i}s_{j})^{\infty}=e$ means that $s_{i}s_{j}$ has infinite order in the group $W$.) Vinberg [130] developed a theory of representations of $W$ as a reflection group in a finite-dimensional real vector space $V$: these are by definition representations $\rho:W\to\mathrm{GL}(V)$ such that each $\rho(s_{i})$ is a linear reflection in a hyperplane of $V$ and the configuration of these reflections is such that $\rho$ is injective, discrete, and the associated fundamental polytope has nonempty interior. These representations may preserve a nondegenerate quadratic form on $V$ (e.g. the image of $\rho$ could be a discrete subgroup of $\mathrm{O}(n,1)$ generated by orthogonal reflections in the faces of a right-angled polyhedron of $\mathbb{H}^{n}$ as in Section 3), but in general they need not preserve any nonzero quadratic form. Representations of $W$ as a reflection group constitute a subset $\mathrm{Hom}_{\mathrm{refl}}(W,\mathrm{GL}(V))$ of $\mathrm{Hom}(W,\mathrm{GL}(V))$ which is semialgebraic (defined by finitely many equalities and inequalities).
@@
Let $W$ be a Coxeter group in $N$ generators as in (7). Suppose $W$ is infinite and Gromov hyperbolic. Then for any $d\geq N$ there exist representations $\rho:W\to\mathrm{SL}^{\pm}(d,\mathbb{R})$ of $W$ as a reflection group which are strongly convex cocompact in $\mathbb{P}(\mathbb{R}^{d})$; for $d\geq 2N-2$, they constitute the full interior of $\mathrm{Hom}_{\mathrm{refl}}(W,\mathrm{GL}(d,\mathbb{R}))$. By Theorem 5.20, these representations are $P_{1}$-Anosov.
@@
One can also use the geometric interpretation of Anosov representations from Theorem 5.20 to prove that free products $\Gamma_{1}*\Gamma_{2}$ of Anosov subgroups $\Gamma_{1},\Gamma_{2}$ are Anosov [55], using a generalisation of the ping pong arguments of Sections 12. For instance, for $11\leq i\leq d-$ 1, let $\tau_{i}:\mathrm{SL}(d,\mathbb{R})\to\mathrm{SL}(S^{2}(\Lambda^{i}\mathbb{R}^{d}))$ be the second symmetric power of the $i$-th exterior power of the standard representation as in Corollary 5.22, let $V^{\prime}_{i}:=S^{2}(\Lambda^{i}\mathbb{R}^{d})\oplus\mathbb{R}$, and let $\tau^{\prime}_{i}:\mathrm{SL}(d,\mathbb{R})\to\mathrm{SL}(V^{\prime}_{i})$ be the direct sum of $\tau_{i}$ and of the trivial representation. Then the following holds.
@@
Finally we note that, although we have seen many constructions of Anosov representations above, it is expected that not every linear Gromov hyperbolic group admits an Anosov representation into some noncompact semisimple Lie group; a concrete example remains to be found.
3 Generalisations of Anosov subgroups
In the past few years, several fruitful generalisations of Anosov subgroups have appeared, which are currently being actively investigated. These generalisations exploit both the dynamical definition of Anosov subgroups from Section 1 and their geometric characterisation from Section 2. Let us briefly mention three of these generalisations.
More general convex cocompact subgroups
We just saw in Theorem 5.20 and Corollaries 5.21 and 5.22 that Anosov representations can be characterised geometrically by a strong convex cocompactness condition in projective space. Here strong refers to the regularity imposed on the properly convex open set $\Omega$ (its boundary $\partial\Omega$ should be $C^{1}$ and contain no segments).
@@ 3 Generalisations of Anosov subgroups
In fact, Weisman [131] has recently given a dynamical characterisation of convex cocompact representations of $\Gamma_{0}$ that extends the characterisation of Anosov representations of Theorem 5.13.(6). The expansion now takes place in various Grassmannians (not only projective space): namely, at each face of the full orbital limit set in $\partial\Omega$, there is expansion in the Grassmannian of $i$-planes of $\mathbb{R}^{d}$ where $i-1$ is the dimension of the face.
@@
As a generalisation of Example 5.25, let $W$ be a Coxeter group in $N$ generators as in (7). Suppose $W$ is infinite. Then there exists a representation $\rho\in\mathrm{Hom}_{\mathrm{refl}}(W,\mathrm{GL}(d,\mathbb{R}))$ which is convex cocompact in $\mathbb{P}(\mathbb{R}^{d})$ for some $d$ if and only if any affine standard subgroup of $W$ is of type $\tilde{A}_{k}$ for some $k\geq 1$ and $W$ does not contain a direct product of two infinite standard subgroups. If this holds, then we can take any $d\geq N$ and the convex cocompact representations then constitute a large open subset of $\mathrm{Hom}_{\mathrm{refl}}(W,\mathrm{GL}(d,\mathbb{R}))$: see [56, § 1.5].
Examples 5.29, 5.30, and 5.31 provide many convex cocompact groups which are not Gromov hyperbolic. (In Example 5.31, the group $W$ is nonhyperbolic as soon as it contains an affine standard subgroup of type $\tilde{A}_{k}$ with $k\geq 2$, see Remark 4.11.)
Some of these groups are still relatively hyperbolic: e.g. in Example 5.31, the group $W$ is relatively hyperbolic with respect to a collection of virtually abelian subgroups of rank $\geq 2$ (see [56, Cor. 1.7]). We refer to [93] for general results about the structure of relatively hyperbolic groups which are convex cocompact in $\mathbb{P}(\mathbb{R}^{d})$ and about the geometry of the associated convex sets. On the other hand, Example 5.29 includes, for $d=d^{\prime}(d^{\prime}+1)/2\geq 6$, discrete subgroups of $\mathrm{PGL}(d,\mathbb{R})$ which divide a symmetric properly convex open set $\Omega_{\mathrm{sym}}\subset\mathbb{P}(\mathbb{R}^{d})\simeq\mathbb{P}(\mathrm{Sym}_{d^{\prime}}(\mathbb{R}))$ and which are isomorphic to cocompact lattices of $\mathrm{PGL}(d^{\prime},\mathbb{R})$, hence not relatively hyperbolic (see Section 3). Further examples of convex cocompact groups which are not relatively hyperbolic can be constructed e.g. using free products inside larger projective spaces: see [55].
Relatively Anosov subgroups
Kapovich–Leeb [97] and Zhu [137, 138] have developed notions of a relatively Anosov representation of a relatively hyperbolic group into a noncompact semisimple Lie group $G$, which generalise the notion of an Anosov representation of a hyperbolic group into $G$ from Section 1. They obtain various characterisations similar to those of Theorem 5.13. The original definition of Anosov representations using flows (Definition 5.1 and Condition 5.4) is recovered in this more general setting by recent work of Zhu and Zimmer [139].
Extending Fact 5.6, if $\Gamma_{0}$ is relatively hyperbolic with respect to a collection of subgroups (called peripheral subgroups), then relatively Anosov representations of $\Gamma_{0}$ into a given $G$ are stable under small deformations that preserve the conjugacy class of the image of each peripheral subgroup [97, 139].
@@
  • some of the Schottky groups of Section 2,
  • the images of certain compositions $\tau\circ\sigma_{0}:\Gamma_{0}\to G$ where $\sigma_{0}:\Gamma_{0}\to G^{\prime}$ is a geometrically finite representation into a semisimple Lie group $G^{\prime}$ with $\mathrm{rank}_{\mathbb{R}}(G^{\prime})=1$ and $\tau:G^{\prime}\to G$ is a representation with compact kernel (e.g. Fact 5.7 generalises to the relative setting);
  • similarly to Section 3, small deformations in $G$ of such $\tau\circ\sigma_{0}(\Gamma_{0})$, preserving the conjugacy class of the image of each peripheral subgroup;
  • certain representations of $\mathrm{PSL}(2,\mathbb{Z})$ into $\mathrm{PGL}(3,\mathbb{R})$ constructed by Schwartz [123] by iterating Pappus’s theorem (see [140, § 13]);
  • for a finite-volume hyperbolic surface $S$, the images of positive (in the sense of Fock–Goncharov [70]) type-preserving representations of $\Gamma_{0}=\pi_{1}(S)$ into a real split simple Lie group $G$, see [43] (for closed $S$, these coincide with the Hitchin representations of Sections 34);
@@
It would be interesting to determine whether relatively Anosov representations of relatively hyperbolic groups can also be fully characterised geometrically similarly to Theorem 5.20 and Corollaries 5.215.22.
Extended geometrically finite subgroups
@@
Recently, Weisman [132, 133] has introduced a notion of extended geometrically finite (or EGF for short) representation of a relatively hyperbolic group. This is a dynamical notion, which extends a dynamical characterisation of Anosov representations in terms of multicones [26]. EGF representations include all Anosov or relatively Anosov representations, all representations of relatively hyperbolic groups which are convex cocompact in the sense of Definition 5.28, as well as other examples (see [133, Th. 1.5–1.7] and [23, Prop. 6.5 & Rem. 6.2]). They are stable under certain small deformations, called peripherally stable, for which the dynamics of the peripheral subgroups does not degenerate too much.
@@
On the other hand, it would be interesting to define a general notion of geometric finiteness in convex projective geometry (involving properly convex open subsets $\Omega$ of $\mathbb{P}(\mathbb{R}^{d})$ where $\partial\Omega$ may contain segments or not be $C^{1}$), and to make the link with Weisman’s EGF representations. A good notion of geometric finiteness should contain as a particular case the notion of convex cocompactness from Definition 5.27. More precisely, a convex projective manifold $M=\Gamma\backslash\Omega$ should be geometrically finite if its convex core $\Gamma\backslash\mathcal{C}^{\mathsf{cor}}_{\Gamma}(\Omega)$ (see Definition 5.27) is covered by a compact piece and finitely many ends of $M$, called cusps, with a controlled geometry. It is not completely clear what the right definition of a cusp should be. Following Cooper, Long, and Tillmann [48], one could define a (full) cusp to be the image in $M$ of some convex open subset of $\Omega$ whose stabiliser in $\Gamma$ is infinite and does not contain any hyperbolic element (i.e. any element of this stabiliser has all its complex eigenvalues of the same modulus); in that case, the cusp is diffeomorphic to the direct product of $\mathbb{R}$ with an affine $(d-2)$-dimensional manifold called the cusp cross-section, and the stabiliser of the cusp is virtually nilpotent [48, Th. 5.3]. The cusp is said to have maximal rank if the cross-section is compact. A more general notion of cusp of maximal rank, where the stabiliser may contain hyperbolic elements but is still assumed to be virtually nilpotent, was studied in [5]. A notion of geometric finiteness involving only such generalised cusps of maximal rank was introduced and studied in [136], where it was characterised in dynamical terms. Examples (both of finite and infinite volume) were constructed in [6, 24] as small deformations of finite-volume real hyperbolic manifolds, using a stability result from [49]; the corresponding representations are EGF by [133]. On the other hand, the study of convex projective cusps of nonmaximal rank, possibly allowing for hyperbolic elements, is still at its infancy, and a good general notion of geometric finiteness in this setting still remains to be found, together with appropriate dynamical characterisations.
@@
Acknowledgements. These notes grew out of a minicourse given at Groups St Andrews 2022 in Newcastle. I would like to warmly thank the organisers (Colin Campbell, Martyn Quick, Edmund Robertson, Colva Roney-Dougal, and David Stewart) for a very stimulating and enjoyable meeting.